0.1 Character Name

* Player Name * Character

* Concept

* Nation

* Religion

* Reputations

* Wealth

0.1.1 Arcana

0.1.2 Stories

* Name * Goal

* Reward

* * Step 1

0.1.3 Traits

* Brawn * Finesse

* Resolve

* Wits

* Panache

0.1.4 Skills

* Aim * Athletics

* Brawl

* Convince

* Empathy

* Hide

* Intimidate

* Notice

* Perform * Ride

* Sailing

* Scholarship

* Tempt

* Theft

* Warfare

* Weaponry

Rank 3 Re-Roll a single die Rank 4 Sets of 15 = 2 Raises Rank 5 10s explode (+1 die)

0.1.5 Death Spiral

* 1 - +1 bonus die to all risks * 2 - Villains gain +2 bonus dice

* 3 - Your 10s explode (+1 die)

* 4 - You become helpless

0.1.6 Backgrounds

0.1.7 Advantages

0.1.8 20 Questions

These are 20 Questions as stated in the 7^{th} Sea Core Rulebook: 2^{nd} Edition.

1. What Nation is your Hero From?

Eisen

2. How would you physically describe your Hero?

Height of 172 cm, with broad shoulders. Female. Blond long hair, in two plaits. Green Eyes. Broad hips.

Heavy cotton dress in forest green, with large iron buttons.

Bloomers, chemise of light cotton worn under dress.

Dark oilskin cloak with a hood.

Iron cross Weis symbolising religious freedom.

Two layer cotton gloves.

3. Does your Hero have recurring mannerisms?

Believes that beauty can be found in everything, and in doing the right thing first.

4. What is your Hero’s motivation?

Believes in personal freedom, provided it does not interfere with others.

0.1.9 Additional Notes

There is no real content here.